a) Graph 𝑔(π‘₯) = π‘₯ 𝑠𝑖𝑛(1⁄π‘₯) to estimate lim 𝑔(π‘₯), zooming in on the origin as necessary π‘₯β†’0

(b) Confirm your estimate in part (a) with a proof.


1
Expert's answer
2021-10-05T11:27:52-0400

(a)





"\lim\limits_{x\to 0}g(x)=\lim\limits_{x\to 0}x\sin(1/x)=0"

(b)

Use the Squeeze Theorem


"-1\leq \sin(1/x)\leq 1, x\in\R, x\not=0"

Then


"-x\leq x\sin(1/x)\leq x, x\in\R, x>0"

"\lim\limits_{x\to 0^+}(-x)=\lim\limits_{x\to 0^+}x=0"

By Squeeze Theorem


"\lim\limits_{x\to 0^+}x\sin(1/x)=0"

"x\leq x\sin(1/x)\leq -x, x\in\R, x<0"

"\lim\limits_{x\to 0^-}(-x)=\lim\limits_{x\to 0^-}x=0"

By Squeeze Theorem


"\lim\limits_{x\to 0^-}x\sin(1/x)=0"

We have


"\lim\limits_{x\to 0^-}x\sin(1/x)=0=\lim\limits_{x\to 0^+}x\sin(1/x)"

Therefore


"\lim\limits_{x\to 0}x\sin(1/x)=0"